OpenStudy (anonymous):

A car traveling initially at +7.7 m/s accelerates uniformly at the rate of +0.76 m/s^2 for a distance of 233m. - what is its velocity at the end of acceleration (answer m/s)

OpenStudy (anonymous):

Known values are v0 = 7.7m/s, a = 0.76 m/s*s, d = 233. The final velocity (v) apparently depends on time (t) and constant acceleration (a) as v = v0 + a*t; (1) From the other side, we know that the distance (d) related with constant acceleration, velocity, and time by an equation d = v0*t+a*t^2/2. (2) So, we can immediately find the t (quadratic equation): t = (-v0+_(v0^2+4ad)^0.5)/a. My calculation gives such answers t1 = 16.62 s, t2 = -36.89 s. The second one seems meaningless, so t = 16,62 s. Substitution that in equation (1) gives v = 20,33 m/s.

OpenStudy (anonymous):

what would it be for 117 m

OpenStudy (anonymous):

Oh, seems I miscalculated above. When distance is 233 m t is 26.32 s and v = 7.7 + 0.76*26.32 = 27.70 m/s. When it is 117 m I got t = 16.67 s and v = 7.7 + 0.76*16.67 = 20.37. That's if I didn't miscalculated again.